# Enhancing Action Recognition by Leveraging the Hierarchical Structure of Actions and Textual Context

### v1

@@ 5.3 Data Fusion
Table 9: Ablation study on the number of attention heads on the fusion transformer.
@@ 5.4 Video Transformer
Input Length: In theory, longer input sequences should improve action recognition accuracy, provided there is no data loss due to excessively long sequences. Table 10 shows the results of our experiments with different sequence lengths. Our findings indicate that an input length of 6.4 seconds (32 input blocks) yields the best performance. Longer sequences tend to miss specific actions within the dataset, while shorter sequences lack the temporal context needed for accurate recognition.
RGB Feature Extractor: Since transformer models for feature extraction are typically trained on static images rather than videos, they may encounter challenges such as motion blur caused by moving objects or people. To address this, we experiment with ViT-H/14 [41] as well as RGB feature extractors from the two-stream network TSN [62], both using pre-trained weights from the Kinetics dataset [14]. Specifically, we evaluate Inception v3 [44] and BN-Inception [65]. As shown in Table 11, despite being trained on static images, the transformer-based method achieves the best accuracy.
Table 10: Ablation study on the number of input blocks.
@@
Table 11: Ablation study on RGB features extractors.
@@
Positional Information: As pointed out in Section 3, extracted frame features lack order information. Table 12 contains an ablation study on using fixed and learnable position encodings against not using them. Results show that positional encoding to preserve the order of the frames is necessary, and that learned positional encoding performs best.
Table 12: Ablation study on position encoding
@@
Number of Encoder Layers and Heads: Finding a balance between the number of encoding layers and attention heads plays a critical role in the performance of transformer models. In Table 14 results show that 4 encoding layers is the best value for this task. Similarly, Table 14 shows that the model accuracy improves when using a single attention head.
Table 13: Ablation study on the number of encoder layers.
@@
Table 14: Ablation study on the number of attention heads.
@@
Embedding Size and Visual Fusion Strategy: We experiment with three different embedding dimensions on two visual fusion strategies: early and late. Early fusion involves concatenating feature vectors before the input to the transformer encoder. On the contrary, a late fusion strategy requires two video encoders, one for RGB features and one for optical flow features. The results are shown in Table 15, with late fusion and an embedding size of 2048 providing the best top-1 accuracy.
Table 15: Ablation study on the embedding dimension and visual fusion strategies.
@@
Encoder Representations: The output of the transformer encoder provides a representation for each input token. To avoid biasing the model towards a particular video block, we experiment with two different representations as explained in Section 3. As shown in Table 16, using a class token improves the encoder’s representation, resulting in a more robust model compared to the mean of each of the encoder’s output tokens.
Table 16: Ablation study on the encoder representation.
@@
6 Conclusion
In this paper, we present a novel approach for action recognition by leveraging the capabilities of language models to generate contextualized representations of previously performed actions and their locations. This enables us to capture the sequential and contextual properties of actions effectively. Additionally, we introduce a joint loss function for both coarse-grained and fine-grained action recognition, benefiting from the hierarchical structure of actions. Our method is built on a transformer encoder for video feature extraction, which incorporates a class token for enhanced representation and positional encoding to preserve the temporal order of input video blocks. We fuse text and video features using a transformer encoder, trained with a dual objective to compute the joint loss.
For our experiments, we extend the TSU dataset by introducing the Hierarchical TSU dataset, which focuses on activities of daily living and emphasizes the hierarchical nature of actions. Despite its simplicity, our method outperforms comparable visual-only methods, including those with the same architecture and pre-trained state-of-the-art models, when trained under the same hyperparameters.
While future work may explore alternative video aggregation and feature fusion mechanisms, this study focuses on the hierarchical properties of actions by analyzing the effects of various textual prompts, such as the number of previous actions and the relevance of location, using different text encoders. We also evaluate different strategies for incorporating hierarchical information into action recognition.
Acknowledgments
This work has been funded by the Valencian regional government CIAICO/2022/132 Consolidated group project AI4Health, and International Center for Aging Research ICAR funded project “IASISTEM”. This work has also been supported by a Spanish national and two regional grants for PhD studies, FPU21/00414, CIACIF/2021/430 and CIACIF/2022/175.
Statements and Declarations
Conflict of interest. The authors declare that they have no conflict of interest.
Data Availability. The dataset used in the current study is publicly available from the corresponding authors. Extended annotations for the Hierarchical TSU dataset are available at https://github.com/3dperceptionlab/HierarchicalActionRecognition.
Code Availability. The code for the presented work has been published on GitHub https://github.com/3dperceptionlab/HierarchicalActionRecognition.

### v2

@@ VI-C Data Fusion
TABLE XIII: Ablation study on the number of attention heads on the fusion transformer.
@@ VI-D Video Transformer
Input Length: In theory, longer input sequences should improve action recognition accuracy, provided there is no data loss due to excessively long sequences. Table XIV shows the results of our experiments with different sequence lengths. Our findings indicate that an input length of 6.4 seconds (32 input blocks) yields the best performance. Longer sequences tend to miss specific actions within the dataset, while shorter sequences lack the temporal context needed for accurate recognition.
RGB Feature Extractor: Since transformer models for feature extraction are typically trained on static images rather than videos, they may encounter challenges such as motion blur caused by moving objects or people. To address this, we experiment with ViT-H/14 [45] as well as RGB feature extractors from the two-stream network TSN [71], both using pre-trained weights from the Kinetics dataset [17]. Specifically, we evaluate Inception v3 [48] and BN-Inception [79]. As shown in Table XV, despite being trained on static images, the transformer-based method achieves the best accuracy.
Positional Information: As pointed out in Section IV, extracted frame features lack order information. Table XVI contains an ablation study on using fixed and learnable position encodings against not using them. Results show that positional encoding to preserve the order of the frames is necessary, and that learned positional encoding performs best.
TABLE XIV: Ablation study on the number of input blocks.
@@
TABLE XV: Ablation study on RGB features extractors.
@@
TABLE XVI: Ablation study on position encoding
@@
Number of Encoder Layers and Heads: Finding a balance between the number of encoding layers and attention heads plays a critical role in the performance of transformer models. In Table XVII results show that 4 encoding layers is the best value for this task. Similarly, Table XVIII shows that the model accuracy improves when using a single attention head.
TABLE XVII: Ablation study on the number of encoder layers.
@@
TABLE XVIII: Ablation study on the number of attention heads.
@@
Embedding Size and Visual Fusion Strategy: We experiment with three different embedding dimensions on two visual fusion strategies: early and late. Early fusion involves concatenating feature vectors before the input to the transformer encoder. In contrast, a late fusion strategy requires two video encoders, one for RGB features and one for optical flow features. The results are shown in Table XIX, with late fusion and an embedding size of 2048 providing the best top-$1$ accuracy.
TABLE XIX: Ablation study on the embedding dimension and visual fusion strategies.
@@
Encoder Representations: The output of the transformer encoder provides a representation for each input token. To avoid biasing the model towards a particular video block, we experiment with two different representations as explained in Section IV. As shown in Table XX, using a class token improves the encoder’s representation, resulting in a more robust model compared to the mean of each of the encoder’s output tokens.
TABLE $X$: Ablation study on the encoder representation.
@@
VII Conclusion
In this work, we proposed a method that effectively leverages both contextual information from past actions, location and domain information, and the hierarchical structure of actions, to improve action recognition accuracy.
To fully leverage these contextual and structural cues, we propose the Hierarchical TSU dataset, an extension of the original TSU with a two-level annotation hierarchy and enriched contextual labels. This dataset supports a more comprehensive evaluation of hierarchical and context-aware action recognition methods.
Through extensive experimentation on the proposed Hierarchical TSU dataset, as well as two additional benchmarks, IkeaASM and Assembly101, we demonstrate that long-term contextual information can be successfully aggregated using language models. In addition, when available, contextual cues such as environment and current location, if available, further contribute to improving action recognition performance.
Hierarchical structures also lead to improvements, although their impact is generally smaller compared to contextual information. Our results highlight the importance of adequately structuring action annotations to fully exploit hierarchical relationships.
Acknowledgment
This work has been funded by the Valencian regional government CIAICO/2022/132 Consolidated group project AI4Health, and by the Spanish State Research Agency (AEI) and ERDF/EU under grant: GEMELIA PID2024-161711OB-I00. This work is also a part of the ENIA Chair of Artificial Intelligence from the University of Alicante (TSI100927-2023-6) funded by the Recovery, Transformation and Resilience Plan from the European Union Next Generation through the Ministry for Digital Transformation and the Civil Service. This work has also been supported by a Spanish national and two regional grants for PhD studies , FPU21/00414, CIACIF/2021/430 and CIACIF/2022/175.